### v1

# Is LLM-as-a-Judge Robust? Investigating Universal Adversarial Attacks on Zero-shot LLM Assessment

Abstract
Large Language Models (LLMs) are powerful zero-shot assessors and are increasingly used in real-world situations such as for written exams or benchmarking systems. Despite this, no existing work has analyzed the vulnerability of judge-LLMs against adversaries attempting to manipulate outputs. This work presents the first study on the adversarial robustness of assessment LLMs, where we search for short universal phrases that when appended to texts can deceive LLMs to provide high assessment scores. Experiments on SummEval and TopicalChat demonstrate that both LLM-scoring and pairwise LLM-comparative assessment are vulnerable to simple concatenation attacks, where in particular LLM-scoring is very susceptible and can yield maximum assessment scores irrespective of the input text quality. Interestingly, such attacks are transferable and phrases learned on smaller open-source LLMs can be applied to larger closed-source models, such as GPT3.5. This highlights the pervasive nature of the adversarial vulnerabilities across different judge-LLM sizes, families and methods. Our findings raise significant concerns on the reliability of LLMs-as-a-judge methods, and underscore the importance of addressing vulnerabilities in LLM assessment methods before deployment in high-stakes real-world scenarios.  [^1] Code: https://github.com/rainavyas/attack-comparative-assessment
Is LLM-as-a-Judge Robust? Investigating Universal Adversarial Attacks on Zero-shot LLM Assessment
Vyas ${}^{\ast}$ University of Cambridge vr313@cam.ac.uk                        Adian ${}^{\ast}$ University of Cambridge al826@cam.ac.uk                        Mark Gales University of Cambridge mjfg@cam.ac.uk
${}^{\ast}$ Equal Contribution.
1 Introduction
Large Language Models (LLMs) have shown to be proficient zero-shot assessors, capable of evaluating texts without requiring any domain-specific training Zheng et al. (2023); Chen et al. (2023); Zhang et al. (2023a). Typical zero-shot approaches prompt powerful LLMs to either generate a single quality score of the assessed text Wang et al. (2023a); Liu et al. (2023b) or to use pairwise comparisons to determine which of two texts are better Liusie et al. (2023); Qin et al. (2023). These zero-shot approaches mark a compelling new paradigm for assessment, enabling straightforward reference-free evaluation that correlates highly with human judgements, while being applicable to a range of diverse attributes. There has consequently been a surge of leveraging LLM-as-a-judge in many applications, including as benchmarks for assessing new models Zheng et al. (2023); Zhu et al. (2023b) or as tools for assessing the written examinations of real candidates.
Figure 1: A simple universal adversarial attack phrase can be concatenated to a candidate response to fool an LLM assessment system into predicting that it is of higher quality. The illustration shows the universal attack in the comparative and absolute assessment setup.
Despite the clear advantages of zero-shot LLM assessment methods, the limitations and robustness of LLM-as-a-judge have been less well-studied. Previous works have demonstrated potential limitations in robustness, and the presence of biases such as positional bias Wang et al. (2023b); Liusie et al. (2023); Zhu et al. (2023b), length bias Koo et al. (2023) and self-preferential behaviours Zheng et al. (2023); Liu et al. (2023d). This paper pushes this paradigm further by investigating whether appending a simple universal phrase to the end of an assessed text could deceive an LLM into predicting high scores regardless of the text’s quality. Such approaches not only pose challenges for model evaluation, where adversaries may manipulate benchmark metrics, but also raise concerns about academic integrity, as students may employ similar tactics to cheat and attain higher scores.
This work is the first to apply adversarial attacks (Szegedy et al., 2014) to zero-shot LLM assessment, and demonstrates that LLM-as-a-judge methods are susceptible to simple concatenative adversarial attacks. Both LLM-scoring and pairwise LLM-comparative assessment are vulnerable to such attacks, and concatenating a universal phrase of under 5 tokens can trick systems into providing significantly higher assessment scores. We find that comparative assessment has mild robustness, but that LLM-scoring is very vulnerable to such attacks and universal attack phrases can cause systems to return a maximum score, irrespective of the input text. We further demonstrate that for LLM-scoring, phrases can be transferable across different model sizes and families, where attack phrases learned on FlanT5-3B can deceive much more powerful systems such as ChatGPT. Finally, as an initial step towards defending against such attacks, we use the perplexity score (Jain et al., 2023) as a simple detection approach, which demonstrates some success. As a whole, our work raises awareness of the vulnerabilities of zero-shot LLM assessment, and highlights that if such systems are to be deployed in critical real-world scenarios, such vulnerabilities should be considered and addressed.
2 Related Work
Bespoke NLG Evaluation.
For Natural Language Generation tasks such as summarization or translation, traditional assessment metrics evaluate generated texts relative to gold standard manual references Lin (2004); Banerjee and Lavie (2005); Zhang et al. (2019). These methods, however, tend to correlate weakly with human assessments. Following work designed automatic evaluation system systems for particular domains and attributes. Examples include systems for dialogue assessment Mehri and Eskenazi (2020), question answering systems for summary consistency Wang et al. (2020); Manakul et al. (2023), boolean answering systems for general summary assessment Zhong et al. (2022a) or neural frameworks for machine translation Rei et al. (2020).
Zero-Shot Assessment with LLMs.
Although suitable for particular domains, these automatic evaluation methods cannot be applied to more general and unseen settings. With the rapidly improving ability of instruction-following LLMs, various works have proposed zero-shot approaches. These include prompting LLMs to provide absolute assessment scores Wang et al. (2023a); Liu et al. (2023b), comparing pairs of texts Liusie et al. (2023); Zheng et al. (2023) or through leveraging assigned output language model probabilities Fu et al. (2023), and in some cases demonstrating state-of-the-art correlations and outperforming performance of bespoke evaluation methods.
Adversarial Attacks on Generative Systems.
Traditionally, NLP attack literature focuses on attacking classification tasks (Alzantot et al., 2018; Garg and Ramakrishnan, 2020; Li et al., 2020; Gao et al., 2018; Wang et al., 2019). However, with the emergence of generative LLMs (Zhao et al., 2023), there has been discussion around NLG adversarial attacks. A range of approaches seek to jailbreak LLMs, and circumvent inherent alignment to generate harmful content (Carlini et al., 2023). Attacks can be categorized as input text perturbation optimization (Zou et al., 2023; Zhu et al., 2024; Lapid et al., 2023); automated adversarial prompt learning (Mehrotra et al., 2023; Liu et al., 2023a; Chao et al., 2023; Jin et al., 2024); human adversarial prompt learning (Wei et al., 2023; Zeng et al., 2024; Liu et al., 2023c); or model configuration manipulation (Huang et al., 2024). Beyond jailbreaking, other works look to extract sensitive data from LLMs (Nasr et al., 2023; Carlini et al., 2020), provoke misclassification (Zhu et al., 2023a) or trick translation systems into making a change in perception (Raina and Gales, 2023; Sadrizadeh et al., 2023). For assessment, although early research has explored attacking NLP assessment systems (Raina et al., 2020), there has been no work on developing attacks for general LLM assessment models such as prompting LLama and GPT, and we are the first to conduct such a study.
3 Zero-shot Assessment with LLMs
As discussed by Zhu et al. (2023b); Liusie et al. (2023), there are two standard reference-free methods of prompting instruction-tuned LLMs for quality assessment:
  • LLM Comparative Assessment where the system uses pairwise comparisons to determine which of two responses are better.
  • LLM Absolute Scoring where an LLM is asked to assign an absolute score to each considered text.
For various assessment methods, we consider rankings tasks where given a query context $\mathbf{d}$ and a set of $N$ responses $\mathbf{x}_{1:N}$, the objective is to determine the quality of each response, $s_{1:N}$. An effective LLM judge should predict scores for each candidate that match the ranking $r_{1:N}$ of the text’s true quality. This section will further discuss the details of both comparative assessment (Section 3.1) and absolute assessment (Section 3.2).
3.1 Comparative Assessment
An LLM prompted for comparative assessment, $\mathcal{F}$, can be used to determine the probability that the first candidate is better than the second. Given the context $\mathbf{d}$ and two candidate responses, $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$, to account for positional bias Liusie et al. (2023); Wang et al. (2023b) one can run comparisons over both orderings and average the probabilities to predict the probability that response $\mathbf{x}_{i}$ is better than response $\mathbf{x}_{j}$,
$p_{ij}=\frac{1}{2}\big{(}\mathcal{F}(\mathbf{x}_{i},\mathbf{x}_{j},\mathbf{d})+(1-\mathcal{F}(\mathbf{x}_{j},\mathbf{x}_{i},\mathbf{d}))\big{)}$ (1)
Note that by doing two inference passes of the model, symmetry is ensured such that $p_{ij}=1-p_{ji}$ for all $i,j\!\in\!\{1,...,N\}$. The average comparative probability for each option $\mathbf{x}_{n}$ can then be used as the predicted quality score $\hat{s}_{n}$,
$\hat{s}_{n}=\hat{s}(\mathbf{x}_{n})=\frac{1}{N}\sum_{j=1}^{N}p_{nj},$ (2)
which can be converted to ranks $\hat{r}_{1:N}$, that can be evaluated against the true ranks $r_{1:N}$.
3.2 Absolute Assessment
In LLM absolute scoring, the LLM, $\mathcal{F}$, is prompted to directly predict an absolute score. The prompt is designed to request the LLM to assess the quality of a text with a score (e.g. between 1-5). Two variants of absolute assessment can be applied; first where the score is directly predicted by the LLM,
$\hat{s}_{n}=\hat{s}(\mathbf{x}_{n})=\mathcal{F}(\mathbf{x}_{n},\mathbf{d}).$ (3)
Alternatively, following G-Eval Liu et al. (2023b), if the output logits are accessible one can estimate the expected score through a fair-average by multiplying each score by its normalized probability,
$\hat{s}_{n}=\hat{s}(\mathbf{x}_{n})=\sum_{k=1:K}kP_{\mathcal{F}}(k|\mathbf{x}_{n},\mathbf{d}),$ (4)
where $K$ is the maximum score, as indicated in the prompt, and the probability for each possible score $k\!\in\!\{1,...,K\}$ is normalized to satisfy basic probability rules, $\sum_{k}P_{\mathcal{F}}(k|\mathbf{x}_{n},\mathbf{c})=1$ and $P_{\mathcal{F}}(k|\mathbf{x}_{n},\mathbf{c})\geq 0$, $\forall n$.
4 Adversarial Assessment Attacks
4.1 Attack Objective
For typical adversarial attacks, an adversary aims to minimally modify the input text ${\mathbf{x}}\rightarrow{\mathbf{x}+\bm{\delta}}$ in an attempt to manipulate the system’s response. The adversarial example ${\bm{\delta}}$ is a small perturbation on the input $\mathbf{x}$, designed to cause a significant change in the output prediction of the system, $\mathcal{F}$,
$\mathcal{F}(\mathbf{x}+\bm{\delta})\neq\mathcal{F}(\mathbf{x}),$ (5)
The small perturbation, $+\bm{\delta}$, is constrained to have a small difference in the input text space, measured by a proxy function of human perception, $\mathcal{G}(\mathbf{x},\mathbf{x}+\bm{\delta})\leq\epsilon$. Our work considers applying simple concatenative attacks to assessment LLMs, where a phrase $\bm{\delta}$ of length $L\!\ll\!|\mathbf{x}|$ is added to the original text $\mathbf{x}$,
$\mathbf{x}+\bm{\delta}=x_{1},\ldots,x_{|\mathbf{x}|},\delta_{1},\ldots,\delta_{L}$ (6)
The attack objective is to then maximally improve the rank of the attacked candidate response with respect to the other candidates. Let $\hat{r}^{\prime}_{i}$ represent the rank of the attacked response, $\mathbf{x}_{i}+\bm{\delta}$, when no other response in $\mathbf{x}_{1:N}$ is perturbed,
$\hat{r}^{\prime}_{i}(\bm{\delta})=\texttt{rank}_{i}\left(\hat{s}(\mathbf{x}_{1}),\ldots,\hat{s}(\mathbf{x}_{i}+\bm{\delta}),\ldots,\hat{s}(\mathbf{x}_{N})\right)$
The adversarial objective is to minimize the predicted rank of candidate $i$ (i.e. the attacked sample) relative to the other unattacked candidates,
$\bm{\delta}^{*}_{i}=\operatorname*{arg\,min}_{\bm{\delta}}(\hat{r}^{\prime}_{i}(\bm{\delta})).$ (7)
In an assessment setting, it is impractical for adversaries to learn an adversarial example $\bm{\delta}^{*}_{i}$ for each candidate response $\mathbf{x}_{i}$. Much more practical is to use a universal adversarial example $\bm{\delta}^{*}$ that could be applied to any candidate’s response $\mathbf{x}_{i}$ to consistently boost the predicted assessment rank. Assuming a training set of $M$ samples of contexts and $N$ candidate responses per context, $\{(\mathbf{d}^{(m)},\mathbf{x}^{(m)}_{1:N})\}_{m=1}^{M}$, the optimal universal adversarial example $\bm{\delta}^{*}$ is the one that most improves the expected rank when attacking each candidate in turn,
$\bar{r}(\bm{\delta})$ $=\frac{1}{NM}\sum_{m}\sum_{n}\hat{r}^{\prime(m)}_{n}(\bm{\delta}).$ (8)
$\bm{\delta}^{*}$ $=\operatorname*{arg\,min}_{\bm{\delta}}(\bar{r}(\bm{\delta}))$ (9)
Where the average is computed over all $M$ contexts and $N$ candidates.
4.2 Practical Attack Approach
In this work, we use a simple greedy search to learn the universal attack phrase  [^2] We also carried out experiments using the Greedy Coordinate Gradient (GCG) attack (Zou et al., 2023) to learn the universal attack phrase, but this approach was found to be not as effective as the greedy search process. Results for GCG experiments are provided in Appendix C.. For a vocabulary, $\mathcal{V}$ the greedy search finds the most effective adversarial word to append iteratively,
$\delta^{*}_{l+1}=\operatorname*{arg\,min}_{\delta\in\mathcal{V}}(\bar{r}(\delta^{*}_{1:l}+\delta)).$ (10)
In practice, it may be computationally too expensive to compute the average rank (as specified in Equation 8). Therefore, we instead approximate the search by greedily finding the token that maximises the expected score when appended to the current sample,
$\delta^{*}_{l+1}$ $=\operatorname*{arg\,max}_{\delta}\mathbbm{E}_{\mathbf{x}}[\hat{s}(\mathbf{x}+\delta^{*}_{1:l}+\delta)]$
Algorithm 1 Greedy Search Universal Attack for Comparative Assessment
$\left\{(\mathbf{c}^{(m)},\mathbf{x}^{(m)}_{1:N})\right\}_{m=1}^{M}$triangleright Training Data
$\mathcal{F}()$triangleright Target Model
$\bm{\delta}^{*}\leftarrow\text{ empty string}$
for $l=1:L$ do
     $a,b\sim\{1,...,N\}$triangleright Select candidate indices
     $\delta^{*}_{l}\leftarrow\text{ none}$
     $q^{*}\leftarrow 0$triangleright Initialize best score
     for $\delta\in\mathcal{V}$ do
         $\bm{\delta}\leftarrow\bm{\delta}^{*}+\delta$triangleright trial attack phrase
         $q\leftarrow 0$
         for $m=1:M$ do
              $s_{a}\leftarrow\mathcal{F}(\mathbf{x}^{(m)}_{a}+\bm{\delta},\mathbf{x}^{(m)}_{b},\mathbf{c}^{(m)})$
              $s_{b}\leftarrow\mathcal{F}(\mathbf{x}^{(m)}_{a},\mathbf{x}^{(m)}_{b}+\bm{\delta},\mathbf{c}^{(m)})$
              $q\leftarrow q+s_{a}+s_{b}$
         end for
         if $q>q^{*}$ then
              $q^{*}\leftarrow q$
              $\delta^{*}_{l}\leftarrow\delta$triangleright Update best attack word
         end if
     end for
     $\bm{\delta}^{*}\leftarrow\bm{\delta}^{*}+\delta^{*}_{l}$triangleright Update attack phrase
end for
Algorithm 2 Greedy Search Universal Attack for Absolute Assessment
$\left\{(\mathbf{c}^{(m)},\mathbf{x}^{(m)}_{1:N})\right\}_{m=1}^{M}$triangleright Training Data
$\mathcal{F}()$triangleright Target Model
$\bm{\delta}^{*}\leftarrow\text{ empty string}$
for $l=1:L$ do
     $a\sim\{1,...,N\}$triangleright Select candidate index
     $\delta^{*}_{l}\leftarrow\text{ none}$
     $q^{*}\leftarrow 0$triangleright Initialize best score
     for $\delta\in\mathcal{V}$ do
         $\bm{\delta}\leftarrow\bm{\delta}^{*}+\delta$triangleright trial attack phrase
         $q\leftarrow 0$
         for $m=1:M$ do
              $s\leftarrow\mathcal{F}(\mathbf{x}^{(m)}_{a}+\bm{\delta},\mathbf{c}^{(m)})$
              $q\leftarrow q+s$
         end for
         if $q>q^{*}$ then
              $q^{*}\leftarrow q$
              $\delta^{*}_{l}\leftarrow\delta$triangleright Update best attack word
         end if
     end for
     $\bm{\delta}^{*}\leftarrow\bm{\delta}^{*}+\delta^{*}_{l}$triangleright Update attack phrase
end for
5 Experimental Setup
5.1 Datasets
We run experiments on two standard language generation evaluation benchmark datasets. The first dataset used is SummEval Fabbri et al. (2021), which is a summary evaluation benchmark of 100 passages, with 16 machine-generated summaries per passage. Each summary is evaluated by human assessors on coherency (COH), consistency (CON), fluency (FLU) and relevance (REL). These attributes can be combined into an overall score (OVE), which is the average of all the individual attributes. The second dataset is TopicalChat Gopalakrishnan et al. (2019), which is a benchmark for dialogue evaluation. There are 60 dialogue contexts, where each context has 6 different machine-generated responses. The responses are assessed by human evaluators on coherency (COH), continuity (CNT), engagingness (ENG), naturalness (NAT), where again the overall score (OVE) can be computed as the average of the individual attributes.
5.2 LLM Assessment Systems
We investigate a range of standard instruction-tuned generative language models in our experiments; FlanT5-xl (3B parameters) Chung et al. (2022); Llama2-7B-chat Touvron et al. (2023); Mistral-7B-chat Jiang et al. (2023) and GPT3.5 (175B parameters). For both comparative and absolute assessment, we perform all attacks on FlanT5-xl. Our prompts for comparative assessment follow the prompts used in Liusie et al. (2023), where different attributes use different adjectives in the prompt. For absolute assessment we follow the prompts of G-Eval Liu et al. (2023b) and use continuous scores (Equation 4) by calculating the expected score over a score range (e.g. 1-5 normalized by their probabilities). Note that the GPT3.5 API does not give token probabilities, and so for GPT3.5 we use standard prompt scoring (Equation 3).
5.3 Methodology
Each dataset is split into a development set and a test set following a 20:80 ratio. We use the development set (20% of the passages) to learn the attack phrase using the simple greedy search to maximise the expected score of the attacked sample (§4.2) and evaluate using the test set (80% of the passages). Furthermore, we only use two of the candidate texts to learn the attacks (i.e. 2 of 16 for SummEval and 2 of 6 for topical chat), and therefore perform the search over a modest 40 summaries for SummEval and 24 responses for TopicalChat.
We learn a single universal attack phrase on the target model, which we keep as FlanT5-xl for all learned phrases. We perform a separate universal concatenation attack for each dataset and attribute, and use the notation ({TASK} {ASSESSMENT} {ATTRIBUTE}) to indicate the task, the assessment type and the evaluation attribute for each learnt universal attack phrase. We specifically consider the following configurations: (SUMM COMP OVE); (TOPIC COMP OVE); (SUMM ABS OVE); (TOPIC ABS OVE); (SUMM COMP CON); (TOPIC COMP CNT); (SUMM ABS CON); (TOPIC ABS CNT)  [^3] The learnt universal attack phrases for each configuration are given in Appendix A.. The specific attributes CON and CNT are selected due to the smallest performance difference in comparative and absolute assessment (Tables 1 and 2). The attack phrases are then assessed on transferability to the other target models; Mistral-7B, Lllama2-7B, GPT3.5. The vocabulary for the greedy attack is sourced from the NLTK python package  [^4] English words corpus is sourced from: nltk.corpus.
5.4 Attack Evaluation
To assess the success of an attack phrase, and for comparing the performance between comparative and absolute, we calculate the average rank of each candidate after an attack is applied (Equation 8). An unsuccessful attack will yield a rank near the average rank, while a very strong attack will provide an average rank of 1 (where each attacked candidate is assumed to be the best of all unattacked candidates of the context).
6 Results
(a) SummEval
(b) TopicalChat
Figure 2: Universal Attack Evaluation (average rank of attacked summary/response) for FlanT5-xl.
6.1 Assessment Performance
Assessment Model OVE COH FLU CON
Comparative FlanT5-xl 54.6 51.2 32.5 47.1
Llama2-7b 31.4 28.2 23.0 27.5
Mistral-7b 25.1 27.6 21.1 27.1
Absolute FlanT5-xl 24.6 27.0 16.6 37.7
Llama2-7b 25.0 28.2 23.0 29.4
Mistral-7b 10.2 14.3 10.5 7.1
GPT3.5 52.5 45.1 38.0 43.2
Table 1: Zero-shot performance (Spearman correlation coefficient) on SummEval. Due to cost GPT3.5 was not evaluated for comparative assessment.
Assessment Model OVE COH CNT ENG
Comparative FlanT5-xl 38.8 47.8 43.5 34.9
Llama2-7b 34.5 35.2 37.1 32.0
Mistral-7b 38.6 33.1 36.1 33.3
GPT3.5 - - - -
Absolute FlanT5-xl 36.2 31.4 43.2 34.9
Llama2-7b 37.1 28.7 20.0 32.9
Mistral-7b 51.7 32.2 37.10 33.5
GPT3.5 56.2 54.7 57.7 49.1
Table 2: Performance (Spearman correlation coefficient) on TopicalChat. Due to cost GPT3.5 was not evaluated for comparative assessment.
Tables 1 and 2 present the assessment ability of each LLM when applied to comparative and absolute assessment for SummEval and TopicalChat. Consistent with literature, comparative assessment performs better than absolute assessment systems for most systems and attributes. However, this comes with greater computational complexity, as $N\!\cdot\!(N-1)$ comparisons are required to compare all pairs of responses (Equation 2), whilst only $N$ inferences are required for absolute assessment. Smaller LLMs (FlanT5-xl, Llama2-7b and Mistral-7b) demonstrate reasonable performance on SummEval and TopicalChat, but larger models (GPT3.5) perform much better, and even when using simple absolute assessment, can outperform smaller models applying comparative assessment.
6.2 Universal Adversarial Attacks
Section 5.3 details the attack configuration and approach used to learn the universal attack phrases. Figure 2 illustrates the impact the universal adversarial attack phrase can have for comparative and absolute assessments on SummEval and TopicalChat evaluation while using FlanT5-xl as the LLM assessment system. For Summeval we attack the overall score (OVE) and the consistency (CON) while for Topical-Chat we attack the overall score (OVE) and the continuity (CNT). The attributes CON and CNT were selected due to the similar observed performance of the absolute and comparative methods on these attributes.
The success of the adversarial attacks is measured by the average ranks of the text after an attack. Figure 2 demonstrates that both comparative assessment and absolute assessment systems have some vulnerability to adversarial attacks, as the average rank decreases, and continues to decrease as more words are added to the attack phrase. However, absolute assessment systems are significantly more susceptible to universal adversarial attacks, and with just four universal attack words, the absolute assessment system will consistently provide a rank of 1 to nearly all input texts. Table 3 provides the raw scores for comparative and absolute assessment, where we see that for absolute assessment, a universal attack phrase of 4 words will yield assessment scores on average near the maximum score of 5. The specific universal attack phrases learnt for each task are given in Appendix A.
Phrase No Attack Attack
SUMM COMP OVE 50.00 51.34
SUMM COMP CON 50.00 57.10
TOPIC COMP OVE 50.00 53.94
TOPIC COMP CNT 50.00 54.06
SUMM ABS OVE 3.73 4.74
SUMM ABS CON 3.88 4.35
TOPIC ABS OVE 2.93 4.63
TOPIC ABS CNT 3.02 4.32
Table 3: Scores for 4-word universal attacks on FlanT5-xl. Note that scores for comparative and absolute assessment are not comparable.
(a) SummEval
(b) TopicalChat
Figure 3: Transferability of universal attack phrases from FlanT5-xl to other models.
The relative robustness of comparative assessment systems over absolute assessment systems can perhaps be explained intuitively. In an absolute assessment setting, an adversary exploits an input space which is not well understood by the model and identifies a region that spuriously encourages the model to predict a high score. However, in comparative assessment, the model is forced to compare the quality of the attacked text to another (unattacked) text, meaning the attack phrase learnt has to be invariant to the text used for comparison. This makes it more challenging to find an effective adversarial attack phrase. Further explanations for the relative robustness of comparative assessment systems are explored in Appendix B.
6.3 Transferability
Figure 2 demonstrated that absolute assessment systems are highly vulnerable to a simple universal attack phrase concatenated to an input text. However, running an adversarial attack on black-box systems may not be feasible due to the large number of calls required for a greedy search, assuming a reasonable vocabulary of $|\mathcal{V}|\approx 50,000$. To circumvent this issue, an adversary can learn the universal attack-phrase on a smaller open-source model (with no detectable limit on the number of model queries) and aim to transfer this attack phrase to other models (e.g., in Demontis et al. (2018)).
The previous universal attack phrases of Figure 2 was learned specifically for FlanT5-xl, which is a relatively poor-performing model. However, Flan-T5 is quite small (3B parameters); one may therefore consider whether attack phrases learned using this smaller model can generalize and deceive larger models. Figure 3 shows that 1) there can be a high level of attack transferability for absolute scoring. For TopicalChat, it appears that the attacks generalize very well to nearly all systems, with all systems being very susceptible to attacks when assessing continuity. 2) It appears that when more powerful models assess the overall (OVE) quality, there can be less effective transferability, suggesting that assessing more general, abstract qualities can be more robust. Further, interestingly it appears that powerful large models (GPT3.5) are more susceptible when attacked by shorter phrases, possibly as the longer phrase may begin to overfit to properties of the attacked model. 3) The attack transfers with mixed success for Summeval, which may highlight that the complexity of dataset can influence attack transferability.
(a) SummEval
(b) TopicalChat
Figure 4: Precision-Recall curve when applying perplexity as a detection defence
6.4 Attack Detection
In this section, we perform an initial investigation into possible defences that could be applied to detect if an adversary is exploiting a system. Defences can take two forms: adversarial training (Goodfellow et al., 2015) where the LLM is re-trained with adversarial examples, or adversarial attack detection where a separate module is designed to identify adversarial inputs. Although recent LLM adversarial training approaches have been proposed (Zhou et al., 2024; Zhang et al., 2023b), re-training is computationally expensive and can harm model performance, hence detection is preferred. Recent detection approaches for NLG adversarial attacks tend to focus on attacks that circumvent LLM safety filters, e.g., generateing malicious content by jailbreaking (Liu et al., 2023c; Zou et al., 2023; Jin et al., 2024). Robey et al. (2023) propose SmoothLLM, where multiple versions of the perturbed input are passed to an LLM and the outputs aggregated. Such defences are inappropriate for LLM-as-a-judge setups, as though the perturbations are designed to cause no semantic change, they can result in changes in other attributes, such as fluency and style, which will impact the LLM assessment. Similarly, Jain et al. (2023); Kumar et al. (2024) propose defence approaches that involve some form of paraphrasing or filtering of the input sequence, which again interferes with the LLM-as-a-judge scores.
A simple and valid defence approach for LLM-as-a-judge is to use perplexity to detect adversarial examples (Jain et al., 2023; Raina et al., 2020). The perplexity is a measure of how unnatural a model, $\theta$ finds a sentence $\mathbf{x}$,
$\texttt{perp}=-\frac{1}{|\mathbf{x}|}\log(P_{\theta}(\mathbf{x})).$ (11)
We use the base Mistral-7B model to compute perplexity. Adversarially attacked samples are expected to be less natural and have higher perplexity. Therefore, we can evaluate the detection performance using precision and recall. We select a specific threshold, $\beta$ to classify an input sample $\mathbf{x}$ as clean or adversarial, where if $\texttt{perp}>\beta$ the sample would be classified as adversarial. The precision, recall and F1 is then
$\texttt{P}=\frac{\text{TP}}{\text{TP+FP}}\quad\texttt{R}=\frac{\text{TP}}{\text{TP+FN}}\quad F1=2\cdot\frac{\texttt{P}\cdot\texttt{R}}{\texttt{P}+\texttt{R}},$
where FP, TP and FN are standard counts for False-Positive, True-Positive and False-Negative respectively. The F1 can be used as a single-value summary of detection performance.
To assess detection, we evaluate on the test split of each dataset, augmented with the universal attack phrase concatenated to each text, such that there is balance between clean and adversarial examples. Figure 4 presents precision-recall (p-r) curves for perplexity detection as the threshold $\beta$ is swept, for the different universal adversarial phrases. Table 4 gives the best F1 scores from the p-r curves. For SummEval all the F1 scores are near 0.7 or significantly above, whilst for TopicalChat the performance is generally even better. This demonstrates that perplexity is fairly effective in disentangling clean and adversarial samples for attacks on LLM-as-a-judge. However, Zhou et al. (2024) argue that defence approaches such as perplexity detection can be circumvented by adaptive adversarial attacks. Hence, though perplexity gives a promising starting point as a defence strategy, future work will explore other more sophisticated detection approaches. Nevertheless, it can also be concluded from the findings in this work that an effective defence against the most threatening adversarial attacks on LLM-as-a-judge is to use comparative assessment over absolute scoring, despite an increased computational cost.
Attack precision recall F1
Summ-CON-2 0.635 0.794 0.706
Summ-CON-4 0.679 0.819 0.742
Summ-OVE-2 0.539 0.988 69.6
Summ-OVE-4 64.7 81.3 72.0
Topic-CNT-2 66.2 84.4 81.7
Topic-CNT-4 74.8 79.5 77.1
Topic-OVE-2 75.2 78.8 76.9
Topic-OVE-4 78.5 85.1 81.7
Table 4: Best F1 (%) (precision, recall) for adversarial sample detection using perplexity. Attack phrases of length 2 words and 4 words considered.
7 Conclusions
This is the first work to examine the adversarial robustness of zero-shot LLM assessment methods against universal concatenation adversarial attacks. Our findings reveal vulnerabilities in both LLM comparative assessment and LLM scoring, with absolute assessment exhibiting particularly high susceptibility. Notably, we demonstrate that attacks devised on smaller systems can effectively transfer to larger ones, underscoring the need for defence strategies. Encouragingly, an initial investigation into detection defences, shows that perplexity can be a promising tool for identifying adversarially manipulated inputs. On the whole, this paper raises awareness around the susceptibility of LLM-as-a-judge NLG assessment systems to universal and transferable adversarial attacks. Further work can explore adaptive attacks and more sophisticated defence approaches to minimize the risk of misuse.
8 Limitations
This paper investigates the vulnerability of LLM-as-a-judge methods in settings where malicious entities may wish to trick systems into returning inflated assessment scores. As the first work on the adversarial robustness of LLM assessment, we used simple attacks (concatenation attack found through a greedy search) which led to simple defences (perplexity). Future work can investigate methods of achieving more subtle attacks, which may require more complex defences to detect. Further, this work focuses on attacking zero-shot assessment methods, however, it is possible to use LLM assessment in few-shot settings, which may be more robust and render attacks less effective. Future work can explore this direction, and also investigate designing prompts that are more robust to attacks.
9 Risks & Ethics
This work reports on the topic of adversarial attacks, where it’s shown that a universal adversarial attack can fool NLG assessment systems into inflating scores of assessed texts. The methods and attacks proposed in this paper do not encourage any harmful content generation and the aim of the work is to raise awareness of the risk of adversarial manipulation for zero-shot NLG assessment. It is possible that highlighting these susceptibilities may inform adversaries of this vulnerability, however, we hope that raising awareness of these risks will encourage the community to further study the robustness of zero-shot LLM assessment methods and reduce the risk of future misuse.
10 Acknowledgements
This work is supported by Cambridge University Press & Assessment (CUP&A), a department of The Chancellor, Masters, and Scholars of the University of Cambridge.
Appendix A Universal Adversarial Phrases
In the main paper, results are presented for a range of universal attack phrases, learnt in different configurations. Further configurations are considered in different sections of the Appendix. For all of these attack phrases, the specific words constituting each phrase are presented in Table 5.
SUMM COMP OVE
E answer E grammatically
SUMM COMP CON
uncontradictory Ay supplementary answer
SUMM ABS OVE
outstandingly superexcellently outstandingly summable
SUMM ABS CON
uncontradictedly undisputably congruity impeccable
TOPIC COMP OVE
informative ending answer E
TOPIC COMP CNT
interester extemporaneous informative answer
TOPIC ABS OVE
informative supercomplete impeccable ovated
TOPIC ABS CNT
continuous superexcellently conformant uncontradictory
SUMM COMP-asymA OVE
E applicableness E E
SUMM COMP-asymB OVE
grammatically sound emendable correctly
SUMM UNI OVE
whoa boggle righto hah
SUMM UNI COH
read inustion newsprint introductorily
SUMM UNI CON
compendent at id id
SUMM UNI FLU
Feuillants cavort extortionately ashore
Table 5: Universal Attack Phrases. Length 1 to length 4 words
Appendix B Analysis of Relative Robustness of Comparative Assessment
It is observed that comparative assessment is more robust than absolute assessment. Arguably this could be due to an implicit prompt ensemble with different output objectives in comparative assessment. In absolute assessment, the adversary has to find a phrase that always pushes the predicted token to the maximal score 5, irrespective of the input test. For comparative assessment, to evaluate the probability summary $i$ is better than $j$ to ensure symmetry, we do two passes through the system. To attack system $i$, for the first pass, the adversary has to ensure the attack phrase increases the probability of token A (the prompt asks the system to select which text input, A or B, is better, where A corresponds to the text in position 1 and B corresponds to the text in position 2) being predicted. For the second pass the adversary has to decrease the predicted probability of token A (as attacked summary is in position 2). This means the objective of the adversary in the different passes is dependent on the prompt ordering of summaries, as well as the objectives being the complete opposite in the two passes (competing objectives). This means the universal attack phrase has to recognise automatically whether it is in position 1 or in position 2 and respectively increase or decrease the output probability of generating token A. This is a lot more challenging and could explain the robustness of comparative assessment. How do we assess this hypothesis:
  • We perform an ablation where the comparative assessment system does asymmetric evaluation such that the probability system $i$ is better than $j$ is measured asymmetrically, with the attacked text always in position 1, such that the adversarial attack only has to maximize the probability of token A. It is expected that the asymmetric comparative assessment system is less robust.
  • We re-apply the greedy search algorithm with this asymmetric setup.
  • We evaluate the efficacy of the attack phrase in the asymmetric setting.
  • We repeat the above experiments with the attack only in position 2 (objective then being to minimize the probability of token B). We term the universal attack phrases asymA and asymB.
The results are presented in Table 6 and Table 7. It seems that even in this asymmetric setting the robustness performance is only slightly (if that) worse than that of the symmetric evaluation setting in the main paper. This suggests that perhaps there is a separate aspect of comparative assessment approach that contributes significantly to the robustness. Further analysis will be required to better understand exactly which aspects of comparative assessment are giving the greatest robustness.
#words s-s s-u u-s u-u all $\bar{r}$
None 45.43 41.07 37.70 42.07 41.54 8.50
1 51.12 51.80 46.68 50.23 50.03 6.17
2 34.96 38.09 34.32 37.54 37.21 9.80
3 48.23 49.04 44.60 47.10 47.06 6.81
Table 6: Direct attack on FlanT5-xl. Evaluating attack phrase SUMM COMP-asymA OVE
#words s-s s-u u-s u-u all $\bar{r}$
None 54.57 62.30 58.93 57.93 58.46 8.50
1 51.91 60.80 52.80 54.36 54.86 9.52
2 57.84 65.04 56.58 58.38 58.90 8.16
3 57.89 63.78 56.29 57.20 57.83 8.54
4 64.70 68.95 60.53 62.00 62.64 7.06
Table 7: Direct attack on FlanT5-xl. Evaluating attack phrase SUMM COMP-asymB OVE
Appendix C Greedy Coordinate Gradient (GCG) Universal Attack
In the main paper we present an iterative greedy search for a universal concatenative attack phrase. Here, we contrast our approach against the Greedy Coordinate Gradient (GCG) adversarial attack approach used by Zou et al. (2023). In our GCG experiments we adopt the default hyperparameter settings from the paper for the universal GCG algorithm. The GCG attack is a whitebox approach that exploits embedding gradients to identify which tokens to substitute from the concatenated phrase. Table 8 shows the impact of incorporating GCG with initialization from the existing learnt attack phrases for absolute assessment and the comparative assessment on overall assessment. From these results it appears that GCG has a negligible impact on the adversarial attack efficacy, and can in many cases degrade the attack (worse average rank) - this is perhaps expected for the best / well optimized attack phrases.
Initialisation No GCG ($\bar{r}$) With GCG ($\bar{r}$)
SUMM COMP OVE 7.96 7.88
SUMM ABS OVE 1.03 2.42
TOPIC COMP OVE 3.16 3.18
TOPIC ABS OVE 1.07 3.56
Table 8: Impact of universal GCG adversarial attack on existing universal attacks
Appendix D Interpretable Attack Results
The main paper presents the impact of the adversarial attack phrases for comparative and absolute assessment systems on the average rank as defined in Equation 8. However, it is more interpretable to understand the the impact on the probability, $p_{ij}$ (Equation 1) of an attacked system being better than other systems for comparative assessment and the impact on the average predicted score (Equation 3) for absolute assessment. Tables 9-12 give the interpretable breakdown of each attack for comparative assessment and Tables 13-28 give the equivalent interpretable breakdown for absolute assessment.
#words s-s s-u u-s u-u $\bar{p}_{ij}$ $\bar{r}$
None 50.00 51.68 48.32 50.00 50.00 8.50
1 50.59 55.97 50.48 52.73 52.80 7.48
2 41.22 49.73 43.90 46.49 46.48 9.75
3 51.27 58.55 51.84 54.33 54.48 6.97
4 50.01 55.88 47.49 51.27 51.34 7.96
Table 9: Direct Attack on FlanT5-xl. Evaluating attack phrase SUMM COMP OVE. SummEval. 16 candidates, with 2 seen candidates (s) and remaining unseen candidates (u).
#words s-s s-u u-s u-u $\bar{p}_{ij}$ $\bar{r}$
None 50.00 53.26 46.74 50.00 50.00 8.50
1 51.65 56.44 48.62 52.04 52.14 7.79
2 52.55 57.70 48.99 52.42 52.62 7.62
3 51.95 56.88 48.38 51.64 51.86 7.93
4 56.64 62.47 53.49 56.85 57.10 6.32
Table 10: Direct Attack on FlanT5-xl. Evaluating attack phrase SUMM COMP CON. SummEval. 16 candidates, with 2 seen candidates (s) and remaining unseen candidates (u).
#words s-s s-u u-s u-u $\bar{p}_{ij}$ $\bar{r}$
None 50.00 44.70 55.30 50.00 50.00 3.50
1 51.25 46.37 56.93 50.13 50.93 3.37
2 55.00 48.11 58.88 52.77 53.34 3.18
3 56.19 49.61 60.14 53.95 54.61 3.06
4 55.18 48.62 59.84 53.33 53.94 3.16
Table 11: Direct Attack on FlanT5-xl. Evaluating attack phrase TOPIC COMP OVE. TopicalChat. 6 candidates, with 2 seen candidates (s) and remaining unseen candidates (u).
#words s-s s-u u-s u-u $\bar{p}_{ij}$ $\bar{r}$
None 50.00 44.27 55.73 50.00 50.00 3.50
1 47.72 44.11 56.19 48.33 49.07 3.55
2 49.81 44.52 56.39 49.04 49.76 3.48
3 53.18 47.88 58.90 52.02 52.76 3.18
4 54.88 48.87 60.07 53.45 54.06 3.12
Table 12: Direct Attack on FlanT5-xl. Evaluating attack phrase TOPIC COMP CNT. TopicalChat. 6 candidates, with 2 seen candidate types (s) and remaining unseen candidates (u).
#words 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 avg $\bar{r}$
None 3.61 3.76 3.79 3.74 3.74 3.76 3.79 3.76 3.65 3.79 3.78 3.77 3.62 3.77 3.67 3.78 3.73 8.50
1 3.96 4.24 4.26 4.19 4.16 4.21 4.19 4.17 3.90 4.20 4.26 4.27 3.99 4.21 4.10 4.24 4.16 2.08
2 4.27 4.49 4.49 4.47 4.44 4.48 4.48 4.41 4.31 4.44 4.48 4.51 4.47 4.47 4.38 4.49 4.44 1.18
3 4.47 4.62 4.63 4.62 4.60 4.63 4.61 4.59 4.46 4.61 4.62 4.64 4.65 4.62 4.56 4.61 4.60 1.07
4 4.70 4.76 4.76 4.75 4.74 4.76 4.75 4.73 4.62 4.74 4.76 4.77 4.75 4.75 4.73 4.75 4.74 1.03
Table 13: Direct Attack on FlanT5-xl. Evaluating attack phrase SUMM ABS OVE. SummEval. 16 candidates.
#words 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 avg $\bar{r}$
None 3.61 3.90 3.94 3.88 3.90 3.93 4.00 3.92 3.74 3.95 3.95 3.96 3.77 3.93 3.74 3.91 3.88 8.50
1 3.83 4.22 4.26 4.18 4.19 4.23 4.19 4.15 3.77 4.17 4.27 4.29 3.98 4.22 3.99 4.21 4.13 3.51
2 3.93 4.27 4.31 4.25 4.25 4.29 4.30 4.23 3.92 4.25 4.32 4.35 4.25 4.27 4.09 4.28 4.22 2.49
3 4.10 4.37 4.38 4.36 4.35 4.39 4.41 4.37 4.25 4.39 4.40 4.42 4.44 4.38 4.24 4.37 4.35 1.71
4 4.10 4.37 4.38 4.36 4.35 4.39 4.41 4.37 4.25 4.39 4.40 4.42 4.44 4.38 4.24 4.37 4.35 1.71
Table 14: Direct Attack on FlanT5-xl. Evaluating attack phrase SUMM ABS CON. SummEval. 16 candidates.
#words 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 avg $\bar{r}$
None 3.00 3.81 3.89 3.75 3.75 3.84 3.88 4.00 3.52 3.96 3.86 3.99 4.00 3.84 3.52 3.52 3.76 8.50
1 3.16 3.80 3.90 3.73 3.73 3.89 3.99 4.00 3.54 3.99 3.91 4.06 3.98 3.80 3.56 3.52 3.78 8.32
2 2.80 3.48 3.59 3.19 3.39 3.41 3.46 3.86 3.01 3.74 3.45 3.52 3.95 3.35 2.99 3.16 3.40 10.47
3 2.80 3.54 3.60 3.24 3.49 3.45 3.61 3.92 2.90 3.74 3.59 3.64 3.99 3.39 3.08 3.21 3.45 10.23
4 3.01 3.64 3.71 3.40 3.51 3.49 3.61 3.98 2.58 3.90 3.61 3.66 3.90 3.50 3.31 3.50 3.52 9.48
Table 15: Transfer Attack on GPT3.5. Evaluating attack phrase SUMM ABS OVE. SummEval. 16 candidates.
#words 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 avg $\bar{r}$
None 3.67 4.05 4.15 4.00 4.00 4.04 4.19 4.05 3.89 4.05 4.12 4.26 4.04 4.01 3.92 3.92 4.02 8.50
1 3.70 4.20 4.24 4.04 4.09 4.26 4.44 4.09 3.91 4.09 4.30 4.61 4.28 4.11 3.94 3.94 4.14 7.63
Table 16: Transfer Attack on GPT3.5. Evaluating attack phrase SUMM ABS CON. SummEval. 16 candidates.
#words 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 avg $\bar{r}$
None 2.08 1.86 1.95 1.83 1.86 1.82 1.87 2.07 1.76 1.99 1.87 1.86 2.04 1.86 1.95 2.09 1.92 8.50
1 2.02 1.89 2.01 1.85 1.90 1.88 1.99 1.98 1.74 1.96 1.95 1.93 1.98 1.87 1.85 2.07 1.93 8.41
2 1.75 1.69 1.80 1.63 1.70 1.68 1.79 1.72 1.63 1.70 1.71 1.76 1.79 1.68 1.63 1.77 1.71 12.38
3 1.73 1.68 1.76 1.65 1.69 1.67 1.75 1.69 1.61 1.70 1.69 1.71 1.81 1.67 1.65 1.75 1.70 12.83
4 1.87 1.79 1.94 1.76 1.81 1.75 1.92 1.85 1.65 1.86 1.81 1.86 1.98 1.79 1.74 1.92 1.83 10.46
Table 17: Transfer Attack on Mistral-7B. Evaluating attack phrase SUMM ABS OVE. SummEval. 16 candidates.
#words 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 avg $\bar{r}$
None 1.64 1.42 1.45 1.46 1.44 1.41 1.40 1.54 1.50 1.51 1.43 1.37 1.47 1.44 1.54 1.57 1.47 8.50
1 1.59 1.44 1.42 1.48 1.45 1.44 1.40 1.53 1.49 1.50 1.42 1.39 1.44 1.46 1.53 1.52 1.47 8.46
2 1.62 1.45 1.41 1.50 1.46 1.46 1.39 1.54 1.55 1.51 1.42 1.38 1.46 1.49 1.56 1.54 1.48 8.02
3 1.52 1.38 1.34 1.41 1.39 1.38 1.33 1.47 1.52 1.45 1.34 1.31 1.38 1.41 1.48 1.45 1.41 10.98
4 1.56 1.40 1.36 1.44 1.42 1.40 1.34 1.50 1.56 1.49 1.37 1.33 1.38 1.44 1.52 1.49 1.44 10.07
Table 18: Transfer Attack on Mistral-7B. Evaluating attack phrase SUMM ABS CON. SummEval. 16 candidates.
#words 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 avg $\bar{r}$
None 3.58 3.74 3.87 3.65 3.72 3.78 3.94 3.73 3.88 3.69 3.80 3.93 3.72 3.70 3.52 3.61 3.74 8.50
1 3.66 3.76 3.87 3.68 3.72 3.76 3.85 3.77 4.02 3.74 3.79 3.86 3.78 3.69 3.56 3.67 3.76 8.31
2 4.23 4.28 4.45 4.26 4.25 4.24 4.33 4.30 4.29 4.28 4.31 4.33 4.21 4.21 4.15 4.24 4.27 3.36
3 4.20 4.23 4.42 4.17 4.21 4.19 4.35 4.28 4.37 4.26 4.24 4.31 4.19 4.18 4.08 4.24 4.24 3.52
4 4.43 4.44 4.58 4.42 4.40 4.39 4.46 4.50 4.41 4.49 4.45 4.43 4.33 4.42 4.35 4.48 4.44 2.30
Table 19: Transfer Attack on Llama-7B. Evaluating attack phrase SUMM ABS OVE. SummEval. 16 candidates.
#words 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 avg $\bar{r}$
None 2.39 2.38 2.38 2.36 2.37 2.39 2.38 2.38 2.27 2.36 2.38 2.38 2.36 2.38 2.37 2.39 2.37 8.50
1 2.38 2.39 2.37 2.38 2.39 2.39 2.37 2.38 2.31 2.37 2.38 2.37 2.39 2.38 2.38 2.40 2.38 8.16
2 2.38 2.39 2.38 2.38 2.39 2.38 2.36 2.38 2.31 2.38 2.37 2.36 2.40 2.39 2.38 2.40 2.38 8.16
3 2.39 2.39 2.37 2.39 2.39 2.38 2.36 2.39 2.36 2.38 2.37 2.36 2.43 2.39 2.40 2.39 2.38 7.81
4 2.40 2.39 2.37 2.39 2.39 2.38 2.36 2.38 2.34 2.38 2.38 2.36 2.41 2.40 2.40 2.39 2.38 7.82
Table 20: Transfer Attack on Llama-7B. Evaluating attack phrase SUMM ABS CON. SummEval. 16 candidates.
#words 1 2 3 4 5 6 avg $\bar{r}$
None 2.98 2.88 2.88 2.88 2.83 3.15 2.93 3.50
1 3.59 3.55 3.59 3.54 3.55 3.85 3.61 1.54
2 4.11 4.13 4.11 4.00 4.03 4.35 4.12 1.22
3 4.44 4.45 4.40 4.33 4.36 4.57 4.42 1.09
4 4.63 4.63 4.61 4.60 4.61 4.67 4.63 1.07
Table 21: Direct Attack on FlanT5-xl. Evaluating attack phrase TOPIC ABS OVE. TopicalChat. 6 candidates.
#words 1 2 3 4 5 6 avg $\bar{r}$
None 3.38 2.54 2.90 2.94 2.67 3.73 3.02 3.50
1 4.92 5.00 4.85 4.88 4.88 4.60 4.85 1.21
2 4.58 4.71 4.90 4.69 4.75 3.96 4.60 1.53
3 4.50 4.77 4.75 4.71 4.48 3.96 4.53 1.61
4 4.35 4.69 4.67 4.69 4.44 3.06 4.32 1.86
Table 22: Direct Attack on FlanT5-xl. Evaluating attack phrase TOPIC ABS CNT. TopicalChat. 6 candidates.
#words 1 2 3 4 5 6 avg $\bar{r}$
None 2.98 2.08 2.42 2.56 2.21 3.19 2.57 3.50
1 3.38 2.88 3.19 3.23 2.90 3.29 3.14 2.64
2 3.23 2.88 3.23 3.44 2.79 3.21 3.13 2.74
3 3.69 3.44 3.94 3.94 3.33 3.35 3.61 2.28
4 2.40 2.46 2.56 2.60 1.83 2.29 2.36 3.79
Table 23: Transfer Attack on GPT3.5. Evaluating attack phrase TOPIC ABS OVE. TopicalChat. 6 candidates.
#words 1 2 3 4 5 6 avg $\bar{r}$
None 3.38 2.54 2.90 2.94 2.67 3.73 3.02 3.50
1 4.92 5.00 4.85 4.88 4.88 4.60 4.85 1.21
2 4.58 4.71 4.90 4.69 4.75 3.96 4.60 1.53
3 4.50 4.77 4.75 4.71 4.48 3.96 4.53 1.61
4 4.35 4.69 4.67 4.69 4.44 3.06 4.32 1.86
Table 24: Transfer Attack on GPT3.5. Evaluating attack phrase TOPIC ABS CNT. TopicalChat. 6 candidates.
#words 1 2 3 4 5 6 avg $\bar{r}$
None 1.63 1.50 1.52 1.51 1.51 1.72 1.57 3.50
1 1.59 1.57 1.59 1.58 1.58 1.70 1.60 3.11
2 1.62 1.58 1.60 1.58 1.58 1.73 1.61 2.98
3 1.59 1.57 1.59 1.58 1.58 1.70 1.60 3.11
4 1.60 1.57 1.61 1.59 1.58 1.73 1.61 2.98
Table 25: Transfer Attack on Mistral-7B. Evaluating attack phrase TOPIC ABS OVE. TopicalChat. 6 candidates.
#words 1 2 3 4 5 6 avg $\bar{r}$
None 2.15 1.85 1.97 2.03 1.81 2.25 2.01 3.50
1 3.33 3.30 3.32 3.27 3.24 3.36 3.30 1.23
2 3.02 3.09 3.17 3.11 3.12 3.25 3.13 1.33
3 3.11 3.10 3.16 3.19 3.15 3.44 3.19 1.26
4 3.23 3.29 3.34 3.28 3.28 3.19 3.27 1.22
Table 26: Transfer Attack on Mistral-7B. Evaluating attack phrase TOPIC ABS CNT. TopicalChat. 6 candidates.
#words 1 2 3 4 5 6 avg $\bar{r}$
None 2.33 2.27 2.31 2.29 2.27 2.46 2.32 3.50
1 2.57 2.66 2.65 2.64 2.67 2.56 2.62 1.57
2 3.28 3.46 3.48 3.47 3.48 3.02 3.37 1.04
3 3.36 3.47 3.49 3.46 3.48 3.15 3.40 1.03
4 3.03 3.13 3.15 3.12 3.12 2.97 3.09 1.09
Table 27: Transfer Attack on Llama-7B. Evaluating attack phrase TOPIC ABS OVE. TopicalChat. 6 candidates.
#words 1 2 3 4 5 6 avg $\bar{r}$
None 2.60 2.58 2.61 2.62 2.59 2.61 2.60 3.50
1 3.28 3.35 3.35 3.34 3.34 3.23 3.31 1.02
2 3.20 3.35 3.40 3.36 3.34 3.06 3.28 1.08
3 3.31 3.50 3.52 3.47 3.46 3.19 3.41 1.03
4 3.11 3.40 3.40 3.36 3.33 3.01 3.27 1.17
Table 28: Transfer Attack on Llama-7B. Evaluating attack phrase TOPIC ABS CNT. TopicalChat. 6 candidates.
Appendix E LLM Prompts
Figure 5 shows the prompts used for absolute scoring via G-EVAL, while Figure 6 shows the prompt template used for comparative assessment.
Figure 5: G-Eval prompt for assessing consistency in Summeval taken from https://github.com/nlpyang/geval. When adapted to TopicalChat, the word ’summary’ is replaced with ’dialogue’ and further minor details are changed for specific attributes
Figure 6: Comparative assessment prompts based on the simple ones used in Liusie et al. (2023). displayed is a prompt for coherency assessment, however different adjectives can be used for different attributes.
Appendix F Attacking Bespoke Assessment Systems
The focus of the paper is on adversarially attacking zero-shot NLG assessment systems. However, one practical defence could be to use a bespoke NLG assessment system that is finetuned to a specific domain. Zhong et al. (2022b) propose such a bespoke system, Unieval that has been finetuned for summary assessment evaluation for each attribute on SummEval. The Unieval system predicts a quality score from 1-5 for each attribute of assessment. Here we explore attacking each attribute of Unieval in turn for the SummEval dataset. Interestingly Unieval appears significantly more robust to these form of adversarial attacks than the zero-shot NLG systems in the main paper. However, it can be observed that there is some vulnerability in the Unieval when assessed on the fluency attribute.
Appendix G Licensing
All datasets used are publicly available. Our implementation utilizes the PyTorch 1.12 framework, an open-source library. We obtained a license from Meta to employ the Llama-7B model via HuggingFace. Additionally, our research is conducted per the licensing agreements of the Mistral-7B, GPT-3.5, and GPT-4 models. We ran our experiments on A100 Nvidia GPU and via OpenAI API.
#words 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 avg $\bar{r}$
None 0.55 0.82 0.80 0.83 0.82 0.86 0.84 0.88 0.61 0.87 0.80 0.90 0.95 0.84 0.76 0.71 0.80 8.50
1 0.55 0.73 0.73 0.73 0.72 0.74 0.73 0.79 0.44 0.79 0.72 0.79 0.71 0.73 0.70 0.68 0.70 12.29
2 0.57 0.76 0.76 0.75 0.75 0.77 0.76 0.82 0.48 0.81 0.75 0.82 0.73 0.76 0.72 0.70 0.73 11.78
3 0.57 0.75 0.76 0.75 0.75 0.77 0.77 0.81 0.49 0.80 0.75 0.83 0.74 0.76 0.71 0.69 0.73 11.80
4 0.57 0.75 0.76 0.74 0.74 0.76 0.77 0.81 0.50 0.80 0.75 0.82 0.72 0.75 0.71 0.69 0.73 11.90
Table 29: Direct Attack on Unieval. Evaluating attack phrase SUMM UNI OVE. SummEval. 16 candidates.
#words 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 avg $\bar{r}$
None 0.38 0.79 0.70 0.83 0.81 0.89 0.86 0.96 0.51 0.95 0.68 0.97 0.97 0.85 0.74 0.58 0.78 8.50
1 0.34 0.61 0.61 0.57 0.60 0.64 0.74 0.76 0.21 0.74 0.58 0.79 0.35 0.62 0.57 0.50 0.58 12.46
2 0.38 0.70 0.66 0.70 0.72 0.77 0.80 0.86 0.29 0.85 0.64 0.86 0.60 0.74 0.69 0.55 0.67 11.77
3 0.35 0.61 0.61 0.57 0.61 0.65 0.73 0.75 0.24 0.74 0.57 0.76 0.41 0.62 0.60 0.50 0.58 12.51
4 0.37 0.63 0.64 0.60 0.64 0.68 0.76 0.77 0.27 0.76 0.60 0.79 0.44 0.64 0.62 0.53 0.61 12.35
Table 30: Direct Attack on Unieval. Evaluating attack phrase SUMM UNI COH. SummEval. 16 candidates.
#words 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 avg $\bar{r}$
None 0.73 0.93 0.94 0.93 0.92 0.94 0.94 0.91 0.58 0.91 0.94 0.95 0.94 0.93 0.86 0.90 0.89 8.50
1 0.77 0.94 0.94 0.94 0.92 0.93 0.93 0.92 0.57 0.92 0.94 0.95 0.94 0.93 0.88 0.91 0.90 8.93
2 0.77 0.94 0.95 0.94 0.92 0.94 0.91 0.92 0.55 0.92 0.95 0.95 0.94 0.94 0.88 0.92 0.90 7.79
3 0.77 0.94 0.94 0.94 0.92 0.94 0.89 0.92 0.57 0.92 0.95 0.95 0.94 0.94 0.88 0.91 0.90 8.27
4 0.77 0.93 0.94 0.93 0.91 0.93 0.90 0.92 0.58 0.92 0.94 0.95 0.94 0.93 0.88 0.91 0.89 9.75
Table 31: Direct Attack on Unieval. Evaluating attack phrase SUMM UNI CON. SummEval. 16 candidates.
#words 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 avg $\bar{r}$
None 0.55 0.75 0.76 0.74 0.72 0.74 0.72 0.77 0.74 0.76 0.77 0.79 0.93 0.74 0.67 0.64 0.74 8.50
1 0.45 0.55 0.57 0.53 0.53 0.54 0.53 0.59 0.40 0.57 0.58 0.60 0.71 0.55 0.51 0.53 0.55 13.21
2 0.62 0.80 0.80 0.80 0.76 0.78 0.71 0.81 0.64 0.80 0.81 0.83 0.92 0.79 0.74 0.70 0.77 7.42
3 0.63 0.80 0.81 0.80 0.77 0.79 0.70 0.81 0.60 0.81 0.82 0.84 0.93 0.80 0.75 0.70 0.77 7.25
4 0.63 0.80 0.81 0.80 0.77 0.79 0.70 0.81 0.60 0.81 0.82 0.84 0.93 0.80 0.75 0.70 0.77 7.26
Table 32: Direct Attack on Unieval. Evaluating attack phrase SUMM UNI FLU. SummEval. 16 candidates.